OpenStudy (anonymous):

Factorization of 2x (to the 2nd power) + x-15

OpenStudy (anonymous):

\[ 2x^{2}+x-15\] what have you tried so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2x^2+x-15 Multiplies to be -30, adds up to be 1. 2x^2+6x-5x-15 (2x^2+6x)(-5x-15) 2x(x+3)-5(x+3) (2x-5)(x+3)=0 x=-3 x=5/2

OpenStudy (anonymous):

Correct thank you

OpenStudy (anonymous):

Factorization of -X^2 +3x +28

OpenStudy (anonymous):

Try it on your own first

OpenStudy (anonymous):

-(x^2-3x-28) Then find the numbers that multiply to be -28 and add to be -3.

OpenStudy (anonymous):

correct, what have you got so far?

OpenStudy (anonymous):

4 and -7 are your factors. So -1(x+4)(x-7)=0